OpenStudy (anonymous):

-Write a system of equations to solve. A farmer has 21 animals on a farm. He only raises sheep and chicken. He counts 62 feet. Dont not solve. Just write the two equations.

OpenStudy (anonymous):

First equation is for Sum of animals: s + c = 21 animals Can you write second equation for sum of legs?

OpenStudy (anonymous):

4s+2c=62 ?? is this right?

OpenStudy (anonymous):

@Chlorophyll

OpenStudy (anonymous):

x=chicken y=sheep so x+y=21 2x+4y=62 SO solve for x and y:)

OpenStudy (anonymous):

completely correct :D

OpenStudy (anonymous):

thank you guys :)